The following income statements were drawn from the annual reports of the Atlanta Company and the Boston Company anta Net sales Cost of goods sold Gross margin Lesss Operating exp 34,600 89,800 16,410)62,380) 27,420 18,190 Selling and admin.exp Net incone 20,880 s 6,230 6,540 All figures are reported in thousands of dollars. Required a-1. Compute the gross margin percentages and return-on-sales ratios of Atlanta and Boston. (Round your answers to the nearest whole number.) a-2. Ascertain which of the company is a high-end retailer based on ratios computed. b. If Atlanta and Boston have equity of $17600 and $21,800, respectively, which company is in the more profitable business? Complete this question by entering your answers in the tabs below Reg A Req A2 Req B Comoute the aross marain percentages and return-on-sales ratios of Atlianta and Boston
media%2F188%2F188b212d-bc5b-41c8-956e-79
media%2F917%2F917354fb-17f4-4597-a579-6a
Answer #1

Atlanta

Boston

Net sales (A)

$34,600

$89,800

Cost of goods sold (B)

$16,410

$62,380

Gross margin (C=A-B)

$18,190

$27,420

Less: Operating expenses

Selling and admin expenses (D)

$11,960

$20,880

Net income (E=C-D)

$6,230

$6,540

1

Ans

Gross margin ratio (Gross margin/Net sales) (C/A)

53%

31%

Return on sales (Net income/Net sales) (E/A)

18%

7%

2

Atlanta company has High end retailer base, since the gross margin and return on sales ratios are higher for Atlanta product

Ans

Atlanta

3

Atlanta

Boston

Equity (A)

17600

21800

Net income (B)

$6,230

$6,540

Return on investment (C=B/A)

35%

30%

Ans

Most profitable business is Atlanta Since higher Return on investment ratio

Atlanta